# Topologically protected flatness in chiral moiré heterostructures

### v1

@@ III.2 Hidden constant of motion
It is, therefore, sufficient to solve for a periodic hidden constant of motion $\vec{X}$ satisfying $[\vec{X}\cdot\vec{\mu},\mathcal{D}]=0$. We rewrite this condition as
$\partial_{\bar{z}}\vec{X}=\alpha M\vec{X},\quad M_{ij}=\varepsilon_{ikj}A_{z}^{(k)},$ (7)
with $\varepsilon_{ijk}$ the Levi-Civita tensor. Note that the matrix $M$ and the vector $\vec{X}$ have complex entries. The fact that this hidden constant of motion is complex is imposed by symmetry. If $\vec{X}$ were real, we could add $(\vec{X}\cdot\mu)\sigma_{3}$ to the Hamiltonian and energetically separate the two flat bands at the magic angle of chiral TBG while preserving all the symmetries of the Hamiltonian, in direct contradiction with the topology of the bands at zero energy [65]. Note also that the antisymmetry of the Levi-Civita tensor in Eq. 7 implies that $x^{2}$ is uniform, while the symmetry of the chiral TBG Hamiltonian ensures that it is real; a demonstration of these facts is given in App. A. Being interested in the small $\alpha$ limit, we look for solutions as a Taylor series $\vec{X}=\sum_{n=0}^{\infty}\alpha^{n}\vec{X}^{(n)}$, whose components must satisfy the recurrence relation $\partial_{\bar{z}}X^{(0)}=0$ and $\partial_{\bar{z}}X^{(n+1)}=MX^{(n)}$ for $n\geq 0$. Solving this recurrence relation yields
$X^{(n)}=I^{(n)}(M)C,\quad I(f)=P\left[\int{\rm d}\bar{z}\,(Mf)\right],$ (8)
where $P[f]=f-{\rm mean}(f)$ removes the uniform part of the periodic function $f$, $I^{(n)}(M)=I\circ I\circ\cdots\circ I\circ I(M)$ denotes $I$ composed $n$-times (with $I^{(0)}$ the identity), while $C$ is an unconstrained integration constant. Numerically, we evaluate the matrix multiplication by $M$ in real space but perform the anti-holomorphic integration in reciprocal (momentum) space, using fast Fourier transforms to switch from one representation to another efficiently.
Figure 2: Maximum of the $n$-th coefficient of the $\alpha$-Taylor series of the constant of motion $\vec{X}$, in logarithmic scale, for different integration constants $C$. The slopes of the linear fit for $n>10$ (dotted lines) provide the constant $p_{C}$ that allows to lower bound the radius of convergence of the Taylor series. These lower bounds are indicated in boxes.
The explicit solution of Eq. 8 allows us to determine for which $\alpha$ the hidden constant of motion exists by calculating the radius of convergence of the series that defines $\vec{X}$. To that end, we compute the $\ell_{\infty}$-norm of the $\vec{X}^{(n)}$ term and find their asymptotic behavior at large $n$. The results are shown in Fig. 2, where we observe that the norms asymptotically grow with $n$ as power laws: $\max|\vec{X}^{(n)}|\sim e^{np_{c}}$ with $p_{C}$ a real coefficient corresponding to the fitted linear slopes in Fig. 2. The radius of convergence $\alpha_{c}$ of the series defining $\vec{X}$ is then lower bounded by $e^{-p_{c}}$. Because the uniform vector $C$ can be arbitrarily chosen, we have repeated the calculation for $C=[1,0,0]$, $C=[0,1,0]$ and $C=[0,0,1]$. In the two first cases, where the $C$ is in-plane, we find an identical radius of convergence $\alpha_{c}^{\parallel}>0.89$; while the third case, where $C$ lies out-of-plane, gives $\alpha_{c}^{\perp}>0.61$. In practice, any uniform vector $C$ will lead to a constant of motion using Eq. 8, which we can therefore construct for all $\alpha<\max(\alpha_{c}^{\parallel},\alpha_{c}^{\perp})$.
We have arrived at one of our main results: for $\alpha<0.89$, it is possible to recast the chiral Hamiltonian of TBG as two decoupled Dirac Hamiltonian feeling opposite abelian magnetic field ±$\pm B^{\prime}$ (see Eq. 6). The effective magnetic field $B^{\prime}$ is represented in Fig. 3a-c for different values of $\alpha$ using our numerical solution for the constant of motion $\vec{X}$. Using graphene’s Fermi velocity and typical values of the inter-layer tunneling and $AB$ sites in TBG, the identified abelian regime corresponds to twist angles $\theta>0.7^{\circ}$, which only includes the first magic angle.
III.3 Topologically protected magic angle
Within the identified abelian regime ($\alpha<0.89$), we can use the AS theorem to predict the emergence of topologically protected flat bands. The only technical subtlety is to identify the minimal compact integration surface $S$ containing all independent fermionic degrees of freedom of the theory to avoid double counting the zero modes [48]. Using the translation and $C_{2y}$ symmetries of TBG, we identify this minimal surface with the half moiré unit cell depicted in Fig. 3c with a dashed triangle.
Figure 3: Magnetic field strength $B^{\prime}=\vec{\nabla}\times\vec{A}^{\prime}$ after abelianization with $C=$[0,0,1] for $\alpha=0.25$ (a), $\alpha=0.4$ (b) and $\alpha=0.55$ (c); the real-space boundaries are the same as in Fig. 1. The minimal integration surface $S$ containing all independent fermionic degrees of freedom of the theory is shown as a dotted triangle in (c). d) Flux of $B^{\prime}$ through $S$ as a function of $\alpha$. The Atiyah-Singer index ensures the existence of a flat band when this flux $\Phi$ reaches $2\pi$. In our calculations, this occurs at $\alpha^{*}=0.581$. We used $512$ points along each primitive vector of the moiré Brillouin zone.
The index theorem (Eq. 2) ensures the existence of a perfectly flat and topologically protected band when the flux $\Phi$ of $B^{\prime}$ through $S$ reaches an integer multiple of $2\pi$ [48]. In Fig. 3d, we show the numerically evaluated value of this flux as a function of $\alpha$. We observe that, although $B^{\prime}$ becomes more and more localized as $\alpha$ grows (Fig. 3a-c), its maximum value also increases and so does its integrated flux. The flux $\Phi$ reaches $2\pi$ for $\alpha^{*}=0.581$, which agrees to less than a percent with the exact value of Ref. [56]. We attribute this difference to the limited numerical accuracy of our calculations as $\alpha$ approaches $\alpha_{c}^{\perp}$. Indeed, the number of terms necessary for the Taylor series to converge increases as $\alpha$ grows, e.g. it reaches $\sim 200$ for $\alpha=0.5$. At this point, the coefficients admixed when numerically solving for the recurrence $I^{(n)}(M)$ differ by more than fifteen orders of magnitudes (which can be seen by extrapolation of Fig. 2 to $n=200$), which corresponds to our numerical accuracy.
Let us finally mention that the counting of flat bands agrees with the one expected for twisted bilayer graphene [40, 41]. Indeed, under the decoupling of Eq. 5, we can split $\mathcal{D}^{\prime}=\mathcal{D}^{\prime}_{+}\otimes\mathcal{D}^{\prime}_{-}$ into two independent abelian Dirac operators feeling opposite gauge field $\mathcal{D}^{\prime}_{\pm}=-2i\partial_{\bar{z}}\pm A^{\prime}_{z}$. As a result, when $\Phi=2\pi$ at $\alpha=\alpha^{*}$, the AS theorem ensures the existence of two topologically protected flat bands. One is fully localized on the $A$ sublattice and is spanned by zero modes of $\mathcal{D}^{\prime}_{+}$, and the other flat band lives on the $B$ sublattice and is spanned by zero modes of $\mathcal{D}^{\prime\dagger}_{-}$. The two sets of zero modes are related to one another by the $C_{2y}$ symmetry of the chiral TBG Hamiltonian.
In essence, the invertible transformation $U$ of Eq. 5 allows us to define layer-spinors that everywhere follow the direction of the fictitious gauge field $\mathcal{A}_{z}$ created by the $AB$ inter-layer tunneling. This decoupling generically relies on an adiabatic approximation, whose region of validity has been analytically identified by explicit computation of the radius of convergence of the small-$\alpha$ Taylor series defining the layer-spinors.
III.4 No anomaly for higher magic angles
We have explicitly constructed an invertible transformation decomposing the chiral model of TBG into two independent abelian components for $\alpha<0.89$ by a Taylor expansion in $\alpha$, enabling us to interpret the flat band at the first magic angle as topologically protected by the AS index theorem. This construction hints that higher magic angles $\alpha=2.221,3.751,\cdots$ are not topologically protected as they lie deep in the non-abelian regime (see Fig. 1). However, it is unclear whether these flat bands are genuinely accidental, i.e. not protected by an index, or if they are topologically protected but the abelianization of the theory involves a transformation $U$ that cannot be expressed as a Taylor series in $\alpha$. Here, we show that the former is true.
Figure 4: Total density in the top layer $\rho_{\rm tot}^{\rm top}(r)$ of the flat band polarized on the $A$ sublattice at the (a) first $\alpha=0.586$, (b) second $\alpha=2.221$, and (c) third $\alpha=3.751$ magic angles. The main difference between (a) and (b-c) is the existence of nodes away from the origin and corners of the Wigner-Seitz moiré unit cell (shown in black) for the higher magic angles.
We rely on the known physical difference between the first and the higher magic angle flat bands of chiral TBG: their nodal structure. The total layer-polarized density of the higher magic angle flat bands is indeed known to possess nodes away from the center and corners of the Wigner-Seitz moiré unit cell ($AA/AB/BA$ points), as shown in Fig. 4 [66]. At these points, the flat-band eigenvectors are fully layer polarized, and the transformation matrix $U$, if it existed, would be entirely carried by $\mu_{3}$. At a nodal point $r_{0}$ different of the center and corners of the moiré Wigner-Seitz cell, we would then have $\vec{X}(r_{0})\cdot\vec{\mu}=x\mu_{3}$, which we show is impossible in App. A if we want to preserve a uniform value of $x$ since both $\vec{A}_{1}$ and $\vec{A}_{2}$ are non-zero at generic point $r_{0}$ of the moiré unit cell. The presence of nodes in the total layer-polarized density of the flat band of the smaller magic angles prevents the abelianization of the theory. It ensures that the higher magic $\alpha$ possesses accidental (unprotected) flat bands. This lack of robustness deep in the non-abelian regime of chiral TBG ($\alpha>1$) agrees with recent mathematical large-$\alpha$ asymptotic results [51] and earlier numerical observations [23].
@@ IV Effect of disorder
In the previous section, we identified two types of flat bands in chiral TBG: the topologically protected ones of the first magic angle ($\alpha<0.89$) and the accidental ones appearing at higher magic angles ($\alpha>0.89$). Here, we study their behaviors in the presence of disorder by building a hierarchy between the various possible disorder channels and analyzing how the flat band responds to them. Our results, summarized in Tab. 1, are finally substantiated by numerical simulations.
channel example abelian ($\alpha<0.89$) non-abelian ($\alpha>0.89$)
off-resonant $AA/BB$-tunneling protected by spectral gap smaller gap (exponentially small in $\alpha$ [67])
chiral homo-strain protected by AS index not protected by an index
resonant, non-chiral $\mu_{1,2}$ AB-tunneling protected by localization near $AA$ weakly protected (less localized)
resonant, non-chiral $\mu_{3}$ hetero-strain not protected weakly protected by nodes near $AA$
Table 1: Effect of the different disorder channels on the flat bands in the abelian ($\alpha<0.89$) and non-abelian ($\alpha>0.89$) regimes. Off-resonant terms connect the flat bands to degrees of freedom above the spectral gap, chiral terms anti-commute with the abelian chiral symmetry of both flat bands, while non-chiral resonant term scatter particles between the two flat bands and are assumed to preserve $C_{3}$ on average (Fig. 5). The second and third columns summarize if the flat bands are protected from the different types of disorder and, if so, by which physical mechanism: either scattering is suppressed by a spectral gap, the AS index, or because the wavefunctions are localized in real space, and the disorder matrix elements are suppressed.

### v2

@@ III.2 Hidden constant of motion
Therefore, it is sufficient to solve for a periodic hidden constant of motion $\vec{X}$ satisfying
$\left[\vec{X}\cdot\vec{\mu},\mathcal{D}\right]=0.$ (6)
We now explicitly construct this constant of motion using the zero-energy states of the symmetry-protected Dirac cones at the $\kappa$ and $\kappa^{\prime}$ corners of TBG’s Brillouin zone. More precisely, regardless of the value of $\alpha$, we are guaranteed two zero modes from the $C_{2}T$ symmetry of the Hamiltonian [43], which are pinned at the Brillouin zone corners in the presence of $C_{3}$ symmetry. Because $\mathcal{D}^{\dagger}({r})=\mathcal{D}^{*}(-{r})$, the zero modes of $\mathcal{D}^{\dagger}$ are related to those of $\mathcal{D}$ by $C_{2}T$ transformation:
$\mathcal{D}|\psi_{\kappa/\kappa^{\prime}}({r})\rangle=0,\quad\mathcal{D}^{\dagger}|\psi^{*}_{\kappa/\kappa^{\prime}}(-{r})\rangle=0$ (7)
with $|\psi_{\kappa,\kappa^{\prime}}(\mathbf{r})\rangle$ a layer-spinor that is fully polarized on the $A$ sublattice. From these definitions, it is clear that
$\vec{X}\cdot\vec{\mu}=|\psi_{\kappa}({r})\rangle\langle\psi^{*}_{\kappa}(-{r})|-|\psi_{\kappa^{\prime}}({r})\rangle\langle\psi^{*}_{\kappa^{\prime}}(-{r})|$ (8)
satisfy the commutation relation Eq. 6. In App. B, we prove that $\vec{X}\cdot\vec{\mu}$ is traceless provided that the relative phase between $\psi_{\kappa}$ and $\psi_{\kappa^{\prime}}$ is properly fixed. We also find that the norm $x$ is uniform in space and equals the Dirac velocity $v_{F}(\alpha)$ around the $\kappa/\kappa^{\prime}$ points for all $\alpha$.
Away from the magic angle where $v_{F}(\alpha)\neq 0$ [42], the invertible transformation in Eq. 4 decouples the chiral TBG Hamiltonian into two independent Dirac Hamiltonian in an effective periodic magnetic field whose total flux through each unit cell needs to vanish to uphold the original translation symmetry of the model. The magic angle defined by $v_{F}(\alpha)=0$ results in a rank-deficient constant of motion and a singular transformation $U$. As we will see below, this singularity physically corresponds to a divergence of the effective abelian magnetic field at a point where $(-2\pi)$ flux quantum is threaded. Well-behaved zero modes of the problem must vanish at this singular point, thereby screening this negative flux quantum. Since the total flux through the moiré unit cell vanishes, these zero modes effectively feel an average of $2\pi$ flux per moiré unit cell, leading to a flat band protected by the AS index theorem. We detail these behaviors below.
III.3 Effective magnetic field
Figure 1: Effective magnetic field $B^{\prime}$ in real space at (a) $\alpha=0.2$, (b) $\alpha=0.4$ and (c) $\alpha=0.55$. Symmetry guarantees that $B^{\prime}(-x,y)=-B^{\prime}(x,y)$. The $B^{\prime}$ field becomes more and more localized as $\alpha$ gets closer to the first magic angle. (d) Wavefunction density $|\psi_{\Uparrow}|^{2}$ and (e) effective magnetic field $B^{\prime}$ along the high-symmetry point contour represented by a dashed line in (c) as a function of $\alpha$ (different colors). (f) Dimensionless flux of $B^{\prime}$ integrated over a small circle of radius $0.5/|\kappa|$ around AB stacking points as a function of $\alpha$ up to the first magic angle. At the first magic angle, (d-f) show that the magnetic field diverges at points where ±$\pm 2\pi$ flux are threaded, and that the wavefunction develops a zero at the BA stacking points where $B^{\prime}$ is infinite and negative. As a result, the system effectively screens one flux quantum and feels an effective positive magnetic flux of $2\pi$ per unit cell.
To give a more intuitive understanding of these statements, let us describe the evolution of the system as $\alpha$ increases from zero to the first magic angle. Fig. 1 shows the magnetic field $B^{\prime}$ defined in Eq. 5 for $\alpha=0.2,0.4,$ and $0.55$. Contrary to the model Eq. 3 that is periodic on a $\sqrt{3}\times\sqrt{3}$-extension of the moiré unit cell (represented with a solid line hexagon in Fig. 1), the magnetic field has the same periodicity as the moiré lattice. This is guaranteed by lattice translational symmetry, which demands that $\mathcal{D}^{\prime}({r}+{a}_{1,2})=\mathcal{D}^{\prime}({r})$ (see App. B). In addition, $C_{2y}T$ symmetry requires that $B^{\prime}(-x,y)=-B^{\prime}(x,y)$ and vanish along the both $y$-axis and the lines related to it by $C_{2x}$ and $C_{3z}$ and translation symmetries. Another consequence of $C_{2y}T$ is that the total flux of $B^{\prime}$ vanishes in each moiré unit cell. For small values of $\alpha$, the field is positive in right-oriented triangles and negative in left-oriented ones. Such a simple structure is, however, very different at larger values of $\alpha$, above the first magic angle, where at each triangle, the effective field fluctuates dramatically, and changes sign rapidly.
Below the first magic angle, as $\alpha$ grows toward it, we can observe that the maximum of $B^{\prime}$ strongly increases and gets confined around the AB and BA stacking points (Fig. 1a-c). Precisely at the first magic angle, $B^{\prime}\to\infty$ at the AB point and diverges with an opposite sign at the BA point. This limit can be seen in Fig. 1d. This has important consequences for the zero-energy modes of $\mathcal{D}^{\prime}$, as we now show. Let us focus on one Abelian component, labelled by $\Uparrow$. The zero mode satisfies the differential equation $(-2i\partial_{\bar{z}}+A_{z}^{\prime})\psi_{\Uparrow}=0$, whose solutions are formally given by
$\psi_{\Uparrow}({r})=f(z)e^{-\rho({r})},\quad\nabla^{2}\rho=B^{\prime},$ (9)
with $f(z)$ a function of $z=x+iy$, which we leave unspecified for the moment. Here $\rho$ is the Poisson potential for the effective magnetic field [62]. When $B^{\prime}$ diverges, $\rho$ diverges with the same sign, yielding a node to the wavefunction $\psi_{\Uparrow}({r})$ at the position where $B^{\prime}\to-\infty$. This is exactly what we observe in Fig. 1d-e through numerically solving the spectrum of $\mathcal{D}^{\prime}$ at momentum $\kappa$.
Integrating the flux in a small circle around the AB point, we see that each of these divergences threads precisely one unit of flux through the system (Fig. 1f). To isolate the effects of these $B^{\prime}$ singularities at AB and BA, we split the magnetic field at magic angles as
$B^{\prime}({r})=2\pi\sum_{t}\left[\delta({r}-{r}_{\rm AB}-{t})-\delta({r}-{r}_{\rm BA}-{t})\right]+\tilde{B}({r}),$ (10)
where $t$ runs over moiré Bravais lattice vectors. The zeros of $\psi_{\Uparrow}$ ($\psi_{\Downarrow}$) at the BA (AB) point effectively screen the negative flux quantum threaded by the magnetic field. As a result, the zero modes at the magic angle feel a non-zero net magnetic flux of $2\pi$ in each moiré unit cell and form a Dirac Landau level protected by the AS theorem.
The same singular behavior of $B^{\prime}$ occurs at the AB and BA stacking points for all higher magic angles, as shown in Fig. 2a-c. For higher magic angles, however, there are multiple singular points in $B^{\prime}$ away from these high-symmetry points. This foreshadows the pathological behaviors of the higher magic angles, which will map onto effective magnetic field with much greater fluctuations than the smooth and well-behaved field obtained at the first magic angles (Fig. 2d-f). At its core, this difference between first and higher magic angle can be traced back to the fact that the $SU(2)$ effective magnetic field =×$\mathcal{B}=\nabla\times\vec{\mathcal{A}}+i\vec{\mathcal{A}}\times\vec{\mathcal{A}}$ is — in the original basis — almost abelian for small $\alpha\lesssim 1$ where the first magic angle lies (because $\vec{\mathcal{A}}\times\vec{\mathcal{A}}\propto\alpha^{2}$ is parametrically smaller than ×$\nabla\times\vec{\mathcal{A}}\propto\alpha$), but it is strongly non-abelian for larger $\alpha\gtrsim 1$ where the higher magic appear. As a result, the basis transformation that abelianizes the Hamiltonian need to strongly fluctuate for higher magic angle, giving rise to more non-uniform and less smooth effective fields.
III.4 Landau level and magnetic field at magic angles
To show the connection to Landau level physics in a more formal manner, we explicitly construct the Landau level basis states using arguments originally developed in Ref. [43], to which we refer for additional details. The decomposition of the magnetic field into a singular and fluctuating part carries over to the Poisson field and yield
$e^{-\rho}=e^{-\tilde{\rho}}\prod_{t}\left|\frac{{r}-{r}_{\rm BA}-{t}}{{r}-{r}_{\rm AB}-{t}}\right|,\quad\nabla^{2}\tilde{\rho}=\tilde{B}.$ (11)
The zeros resulting from the singular part of the magnetic field allows to choose the meromorphic function
$f_{k}(z)=e^{2\pi k_{1}z}\dfrac{\theta_{1}(z-z_{\rm BA}-k_{x}-ik_{y}|\omega)}{\theta_{1}(z-z_{\rm BA}|\omega)},$ (12)
in Eq. 9 to build a zero mode $\psi_{\Uparrow,k}$ of $\mathcal{D}^{\prime}$ with a well-defined momentum $k$. Here, $\theta_{1}$ represents the Jacobi theta function, and we have fixed the moiré lattice constant to one, and defined $b_{1,2}=\kappa_{1,2}-\kappa_{0}$ with which we have decomposed $k=k_{1}b_{1}+k_{2}b_{2}$.
One could guess that a similar construction for the second abelian component of $\mathcal{D}^{\prime}$ would give another band $\psi_{\Downarrow,k}$ of zero modes. However, it is not the case. The rank deficiency of $\vec{X}\cdot\vec{\mu}$ at the magic angles ensures that $|\psi_{A,k}\rangle=U[\psi_{\Uparrow,k},0]^{T}$ and $|\psi_{A,k}^{\prime}\rangle=U[0,\psi_{\Downarrow,k}]^{T}$ are linear dependent once transformed back into the original layer-basis of the Hamiltonian. Here, $\psi_{\Uparrow,k}$ denotes the zero-energy wavefunction of the top abelian component of the decoupled Dirac operator $\mathcal{D}^{\prime}$ that carries quasi-momentum $k$. In other words, we recover the desired flatband counting for chiral TBG, with one exact flat band per spin, valley, and sub-lattice.
Figure 2: The effective magnetic field $B^{\prime}$ at the (a) first, (b) second and (c) third magic angles. At all magic angles, $\delta$-function peaks are observed at the AB/BA stacking points, each carrying a flux of $2\pi$ in magnitude. Due to symmetry, the total flux within each unit cell remains zero. Panels (d)-(f) show the effective magnetic field $B^{\prime\prime}$ after applying the singular transformation (Eq. 13) for the first to third magic angles. The total flux in each unit cell now sums to $2\pi$, resulting in the formation of a flat band with an AS index.
To remove the singular component of the magnetic field and obtain an intuitive picture of the magnetic field effectively felt by $|\psi_{A,k}\rangle$, we include the singular part of the magnetic field leading to a node in the eigenstates’ real-space density as follows. Instead of the almost everywhere invertible transformation $U$, we use a singular transformation $\tilde{U}=U\Lambda(r)\mu_{3}$ to define
$\mathcal{D}^{\prime\prime}$ $=\tilde{U}^{-1}\mathcal{D}\tilde{U}=-2i\mu_{0}\partial_{\bar{z}}+A_{z}^{\prime\prime}\mu_{3},$ (13)
$A_{z}^{\prime\prime}$ $=A_{z}^{\prime}+2i\partial_{\bar{z}}\ln\Lambda,\quad B^{\prime\prime}=\nabla\times A^{\prime\prime},$
where choosing
$\Lambda({r})=\frac{\theta_{1}(z-z_{\rm BA}|\omega)}{||U[1,0]^{T}||},$ (14)
eliminates all singularities in the density of the eigenvectors. We compare the original $B^{\prime}$ field with the non-singular and well-behaved $B^{\prime\prime}$ resulting from the singular transformation in Fig. 2, which were first discussed in Ref. [68] for the first magic angle. Note that this singular transformation can only be carried out at the magic angle, where the $B^{\prime}$ field is itself singular, such that the zero-mode wavefunction have nodes at the BA stacking points.
The additional divergences in $B^{\prime}$ of higher magic angle manifest in $B^{\prime\prime}$ as oscillations of larger and larger amplitudes (note the scales in Fig. 2d-f). Though all magic angles admit a Landau level description, we now argue and numerically demonstrate that these strong fluctuations prevent the protection from the index theorem to play any role in realistic settings.
@@ IV Effect of disorder
In the previous section, we identified a hidden constant of motion decoupling the flat bands in chiral TBG, endowing them with an analytical index. Here, we phenomenologically argue how the flat bands of chiral TBG respond to various forms of disorder. Our analysis unveils that only the first magic angle successfully harnesses the topological protection offered by the index theorem. To see this, we classify the disorder channels from the most to the least detrimental for the flat bands in Sec. IV.1. Our results, summarized in Tab. 1, are then substantiated by numerical simulations both in (Sec. IV.2) and away from (Sec. IV.3) the chiral limit.
disorder channel first magic angle higher magic angles example
chiral $\sigma_{1/2}\mu_{0}$ protected protected homo-strain
chiral $P(\vec{X})$
suppressed by $P(X)$ acting locally
not protected hetero-strain
resonant non-chiral
suppressed by wavefunction localization in $k$
not protected $AB$-tunneling
off-resonant suppressed by spectral gap
not protected, gap squeezing
$AA/BB$-tunneling
Table 1: Effect of the different disorder channels on the flat bands. The first two lines describe disorder that anticommute with the chiral symmetry in the decoupled abelian basis of the Hamiltonian, to which the flat bands are perfectly immune thanks to the AS theorem. We split the chiral disorder channels into those independent of the magic angle – the $\sigma_{1/2}\mu_{0}$ homo-strain of the first line – and the fine-tuned channel $P(\vec{X})$ that explicitly depend on the constant of motion $\vec{X}$ at each magic angle. This fine-tuned disorder only has a large overlap with physical disorder realizations that carry no momentum on average at the first magic angle where $P(\vec{X})$ acts locally (see Fig. 3). The third line describes resonant disorder that directly couples the two sublattice polarized flat bands, and encompasses variations of the $AB$-tunneling amplitude. Only the first magic angle shows strong protection against this type of disorder (see also the numerical results of Sec. IV.2) due to the strong localization of $\vec{X}$ in momentum space that mitigates the effects of all the disorder’s high-harmonics (Fig. 3). Finally, the last line concerns off-resonant terms connecting the flat bands to degrees of freedom above the spectral gap. The strong fluctuations of the effective magnetic field at higher magic angle (Fig. 2) yield an exponentially small gap for the higher magic angles, as first proved in Ref. [69] based on generic argument for non-uniform Landau problems.